Convert 1000 µg mm ns-1to kg m s-1


1
Expert's answer
2021-03-21T11:31:25-0400
"1000\frac{\mu g mm}{ns}=1000\frac{10^{-6} kg 10^{-3}m}{10^{-9}s}\\
1000\frac{\mu g mm}{ns}=1000\frac{ kg m}{s}"